a substance in a chemical reaction is used up to a rate which is proportional to the amount of the substance present at any time. if 80% of the substance is used up in 4 hours, how much if the substance is left at the end of 120 minutes
1
Expert's answer
2020-10-18T17:18:30-0400

Let "A" be the amount of the substance present at time "t" .

The rate at which the substance is used up is "-\frac{dA}{dt}" . The negative sign is due to the fact that the amount of substance is decreasing with time.

According to the problem, the rate "-\frac{dA}{dt}" is proportional to "A" .

"\therefore -\frac{dA}{dt}\propto A\\
\Rightarrow -\frac{dA}{dt}=kA"

where "k" is the proportionality constant.

"\Rightarrow \frac{dA}{A}=-kdt"

Let the initial amount of the substance is "A_0" .

Integrating both sides,

"\int_{A_0}^A\frac{dA}{A}=-k\int_0^tdt\\
\Rightarrow\left[ \ln A\right]_{A_0}^A=-k[t]_0^t\\
\Rightarrow \ln A-\ln A_0=-kt\\
\Rightarrow \ln \frac{A}{A_0}=-kt\\
\Rightarrow \frac{A}{A_0}=e^{-kt}\qquad ...............(1)"

At "t=4\ hours" , 80% of the substance is used up and hence 20% of the substance is left.

Therefore, at "t=4\ hours" , "A=20\%\ of\ A_0=0.20A_0"

Substituting these values in Eq.(1),

"\frac{0.20A_0}{A_0}=e^{-4k}\\
\Rightarrow \ln (0.20)=-4k\\
\Rightarrow k=-\frac{1}{4}\ln (0.20)=0.402\ hour^{-1}"

At the end of 120 minutes, i.e, at "t=2\ hours" , from Eq.(1),

"\frac{A}{A_0}=e^{-0.402\times 2}\\
\Rightarrow \frac{A}{A_0}=0.478\\
\Rightarrow A=0.478A_0\\
A=47.8\%\ of A_0"


Answer: At the end of 120 minutes 47.8% of the substance is left.